OpenStudy (anonymous):

3y + 12 / 8y3 divided by 9y + 36 / 15y3

OpenStudy (anonymous):

You need to follow the steps: 1) Organize and simplify: \[(\frac{ 3y+12 }{ 8y^3})(\frac{ 15y^3 }{ 9y+36 })=(\frac{ 3(y+4) }{ 8y^3})(\frac{ 15y^3 }{ 9(y+4) })\] 2) Eliminate (i.e. simplify more): \[(\frac{ 3y+12 }{ 8y^3})(\frac{ 15y^3 }{ 9y+36 })=(\frac{ 3(y+4) }{ 8y^3})(\frac{ 15y^3 }{ 9(y+4) })=(\frac{ (y+4) }{ 8y^3})(\frac{ 5y^3 }{ (y+4) })=\frac{ (y+4) }{ 8})(\frac{ 5 }{ (y+4) })\] I will leave the rest to you.

OpenStudy (anonymous):

Thanks!!!!!!!!